OpenStudy (griffindore):

Can someone please help me? A closed container has 3.06 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 0.47 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 ⋅ 10–1 0.51 grams, because (3.06 ⋅ 1.67) ⋅(1023 ⋅ 10–24) = 5.1102 ⋅ 10–1 4.73 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 5.11 grams, because (3.06 ⋅ 1.67) ⋅ (1023 ⋅ 10–24) = 5.1102

OpenStudy (whpalmer4):

To multiply two numbers written in scientific notation, you multiply the mantissa portions (the number without an exponent) to get the new mantissa, then add the exponents to get the new exponent.

OpenStudy (griffindore):

i dont really get it?

OpenStudy (whpalmer4):

\[a*10^b * c*10^d = (a*c) * 10^{b+d}\]

OpenStudy (griffindore):

oh so like...

OpenStudy (griffindore):

3.06 x 1.67 x 10^23 x (-24)

OpenStudy (griffindore):

is that right? ^

OpenStudy (whpalmer4):

so far so good. What do you get when you do that?

OpenStudy (griffindore):

give me a sec

OpenStudy (griffindore):

it would be 5.11 x 10^-552

OpenStudy (griffindore):

oh so it would be D

OpenStudy (griffindore):

thx

OpenStudy (whpalmer4):

Hold on...D is NOT correct... \[3.06*10^{23} * 1.67*10^{-24} = 3.06*1.67*10^{23}*10^{-24} = 5.11*10^{23+(-24)} \]\[= 5.11*10^{-1} = 0.511\]